# Recent Advances in Near-Field Beam Training and Channel Estimation for XL-MIMO Systems

### v1

@@ I Introduction
With the commercial rollout of fifth-generation (5G) networks, both academia and industry have been actively exploring key technologies for the sixth-generation (6G) wireless systems. Among these, extremely large-scale multiple-input multiple-output (XL-MIMO) has emerged as a promising candidate. By equipping base stations (BSs) with an order of magnitude more antennas than conventional massive MIMO, XL-MIMO enables significantly higher spatial resolution and spectral efficiency. However, the larger array size, coupled with shrinking cell coverage in modern communication networks, increases the likelihood of user equipments operating in the near-field region. In this scenario, the conventional planar wave assumption becomes inaccurate, necessitating a shift to a more precise near-field spherical wave model. Unfortunately, the shift from planar to spherical wave propagation invalidates many fundamental techniques used in current MIMO systems, including beam training, channel estimation, and beamforming. In particular, the near-field spherical wave model poses significant challenges for beam training and channel estimation, which are critical for realizing the promised gains in spectral efficiency. Thus, investigating their impact and developing effective solutions is essential for the successful deployment of XL-MIMO.
@@
The channel information obtained from beam training is inherently limited, as only the signal strength is fed back. To fully exploit the potential of XL-MIMO and enhance multi-user communications, comprehensive channel estimation is necessary to acquire the complete channel state information (CSI). Conventional channel estimation techniques employed in massive MIMO, such as minimum mean square error (MMSE) and least squares (LS), can be directly extended to time division duplex (TDD)-based XL-MIMO systems with fully digital precoding. However, for TDD systems with hybrid precoding—commonly adopted in XL-MIMO to minimize the number of radio-frequency (RF) chains—and frequency division duplex (FDD) systems, significant challenges arise when applying existing compressive sensing (CS)-based algorithms developed for conventional massive MIMO. These methods typically exploit channel sparsity in the angular domain to reduce pilot overhead, an assumption that no longer holds under the near-field spherical wave propagation model. Hence, novel channel estimation algorithms that jointly leverage sparsity in both the angular and distance domains are essential for effective CSI acquisition in near-field XL-MIMO systems.
Figure 1: Near-field channel: a) spherical wave illustration; b) beam energy-spread effect (received beam pattern under far-field beamformers for different user angles and ranges under 256 antennas at 30 GHz); c) finite depth of beamfocusing (512 antennas at 100 GHz).
@@
This article is organized as follows: Section II introduces the fundamentals of XL-MIMO. Sections III and IV present the key techniques for beam training and channel estimation in XL-MIMO systems, respectively. Section IV highlights potential challenges and open directions, while Section V draws the conclusion.
Figure 2: Beam training: a) polar-domain-based beam training [1]; b) effective Rayleigh distance [2].
@@ II Fundamentals of XL-MIMO Systems
XL-MIMO has garnered significant attention as a key enabler for 6G and beyond networks. By deploying an order of magnitude more antennas than conventional massive MIMO at the BSs, XL-MIMO not only enhances multi-user connectivity but also substantially improves spectral and energy efficiency. Notably, studies have demonstrated that XL-MIMO can achieve a considerable capacity gain over existing MIMO systems. However, the dramatic increase in array size leads to an extended Rayleigh/Fraunhofer distance, a fundamental criterion for delineating the near-field and far-field regions. Concurrently, the continuous reduction in cell size in modern communication networks exacerbates this effect, making it increasingly unlikely for user equipment to remain within the conventional far-field region. In the near-field regime, the conventional planar wave assumption becomes invalid, necessitating the adoption of a more accurate spherical wave-based channel model (shown in Fig. 1(a)) to capture the spatial characteristics of signal propagation.
@@
In a spherical wave-based channel model, the phase of the array response vector is no longer a linear function of the array element indices, as observed in the conventional planar wave assumption. Instead, it is intricately coupled with the element-wise propagation distance in a nonlinear manner. This nonlinear phase-distance relationship significantly alters the spatial characteristics of wavefront propagation in XL-MIMO systems. As illustrated in Fig. 1(b), this phenomenon gives rise to the so-called beam energy-spread effect when applying conventional DFT-based beamforming [1]. Specifically, in the far-field regime, DFT beams are designed under the assumption of a linear phase progression across array elements, ensuring coherent beamforming gain at the intended direction. However, the phase evolution is nonlinear in the near-field region such that the direct application of DFT beams leads to energy dispersion over a wider angular span, reducing beamforming gain and impairing both beam training and channel estimation accuracy. This degradation poses a fundamental challenge in XL-MIMO system design.
@@
On the other hand, the inherent coupling of phase and distance in the array response vector enables finite-depth beam focusing, a unique property of near-field spherical wavefronts, as shown in Fig. 1(c). Unlike far-field beams, which form a plane wavefront extending infinitely, near-field beams can concentrate energy at a specific spatial location, not only in angular dimensions but also in range. This capability allows BSs to distinguish multiple users that share the same angular direction but are located at different distances—a functionality absent in conventional MIMO systems, where angular separation is the sole basis for spatial multiplexing. The ability to focus beams at precise locations holds great potential for improving user separation, enhancing spatial multiplexing gains, and enabling new applications such as location-aware communications and holographic radio.
@@ III-A Motivation
Although the DFT-based codebook is highly effective for beam training in far-field systems, it is not directly applicable to near-field scenarios. When a DFT-based codebook is used in the near-field, it results in beam energy dispersion, as illustrated in Fig. 1(b), which complicates the identification of the optimal beam angle. Additionally, applying such codewords for beamforming introduces significant power leakage to users with similar angular positions. Furthermore, the DFT-based codebook does not leverage the finite-depth beam focusing capability of the near-field, rendering it incapable of distinguishing users with identical angular coordinates but different radial distances. Therefore, for near-field XL-MIMO systems, novel beam training techniques and codebooks that jointly consider both angular and range dimensions must be developed.
Figure 3: Beam training [3]: a) far-field for angle estimation; b) joint angle and distance estimation.
@@ III-B1 Polar-domain-based beam training
Polar-domain-based beam training has emerged as a promising approach for near-field XL-MIMO systems [1]. As illustrated in Fig. 2(a), this method partitions the entire near-field region into discrete grids, each characterized by distinct angular and range coordinates. The codebook is then designed such that each codeword generates a beam that concentrates the majority of its energy within a single grid, effectively incorporating both angular and range information into the beam training process. Moreover, to minimize power leakage into adjacent beams, uniform sampling should be applied to the angular dimension, while non-uniform sampling is required for the range dimension to better capture the characteristics of near-field propagation.
@@
[…]
Figure 4: Beam training [4]: a) multi-beam training; b) remove angle ambiguity.
@@ III-B2 Conventional DFT-based beam training
Due to the beam energy spread, conventional DFT-based codebooks cannot be directly applied for angle estimation in the near-field. However, recent studies have demonstrated that the received beam pattern from a conventional DFT-based codebook can still be leveraged for user angle estimation [5, 6, 7]. In [5], a two-phase near-field beam training method was proposed to sequentially estimate the user’s angle and range, thereby mitigating the need for a computationally expensive two-dimensional exhaustive search. Specifically, the conventional DFT-based codebook is initially employed for beam sweeping. As illustrated in Fig. 1(b), the true user angle is approximately centered within the angular support—defined as the region where the normalized beam gain exceeds a given threshold. Once the user angle is determined, the user range is estimated via a one-dimensional exhaustive search utilizing the polar-domain codebook. Notably, the beam sweeping in [5] employs the entire antenna array, which can lead to excessive training overhead, especially in XL-MIMO systems where the number of antennas is substantial.
@@
It is important to note that both [5] and [6] utilize the polar-domain codebook for range estimation. As an on-grid beam training method, it only optimally aligns with users whose range coordinates coincide with predefined sampling points. However, since actual user positions are continuous and randomly distributed, any misalignment between the user’s true position and the nearest sampling grid results in beam training loss, thereby degrading overall system performance. A straightforward approach to improve beam training accuracy is to increase the density of range sampling points. However, this comes at the cost of increased training overhead. To address this limitation, an off-grid beam training method was proposed in [7], which eliminates the need for a polar-domain codebook by relying solely on a conventional DFT-based codebook for joint angle and range estimation. In this method, the user angle is first identified using the DFT-based codebook, as in [5]. The user range is then inferred based on the observation that the angular support width at a given user angle decreases as the user range increases, as depicted in Fig. 1(b) (blue and red solid curves). Since the exact functional relationship between angular support width and user range is difficult to derive analytically, a surrogate function is employed to estimate the user range in a semi-closed form.
@@ IV-B1 Channel Estimation in Uplink TDD Systems with Hybrid Beamforming Architecture
Figure 5: Systems with hybrid-field channels.
@@ V-A Real Measurement Data-Based Performance Evaluation
In most existing studies, the performance of beam training and channel estimation is evaluated using artificially generated channels based on the near-field channel model. However, in practical scenarios, real-world channel characteristics may deviate from those predicted by the assumed near-field model due to factors such as hardware impairments, environmental variations, and multipath effects. Consequently, the effectiveness of the proposed beam training and channel estimation techniques under real-world conditions remains uncertain. To bridge this gap, it is crucial to assess their performance using real-world measured channel data, ensuring their robustness and practicality in actual deployment scenarios.
@@ V-C Beam Training and Channel Estimation for FR3 Band
Frequency Range 3 (FR3), spanning approximately 7–24 GHz, is anticipated to be one of the most critical new spectrum bands for 6G. Its propagation characteristics differ significantly from those observed in both mmWave)/THz bands and sub-6 GHz frequencies. However, research on beam training and channel estimation for the FR3 band remains in its early stages, requiring substantial investigation. A key challenge in this context is determining whether FR3 channels will primarily exhibit near-field characteristics or if they will operate in a hybrid regime, incorporating both near-field and far-field components. Additionally, given the broad frequency range of FR3, its lower and upper portions may exhibit distinct propagation behaviors, further complicating channel modeling and system design. Therefore, a deeper understanding of FR3 channel spatial properties and their impact on beam training and channel estimation is essential for optimizing future 6G communication systems.
@@ VI Conclusion
XL-MIMO is poised to play a pivotal role in next-generation wireless communication systems, offering unprecedented improvements in spectral efficiency and spatial resolution. However, the transition from conventional massive MIMO to XL-MIMO introduces new challenges, particularly in beam training and channel estimation, due to the shift from planar to spherical wave propagation. In this article, we provided a structured and in-depth review of the state-of-the-art beam training and channel estimation methods, categorizing existing approaches and analyzing their effectiveness in near-field propagation environments. Looking ahead, future research efforts should focus on developing adaptive and scalable solutions, potentially leveraging sensing-aided techniques to enhance system performance. Moreover, experimental validation using real-world channel measurements will be essential to ensure practical applicability.

### v2

@@ I Introduction
With the commercial rollout of fifth-generation (5G) networks, both academia and industry have been actively exploring key technologies for the sixth-generation (6G) wireless systems. Among these, extremely large-scale multiple-input multiple-output (XL-MIMO) has emerged as a promising candidate [15, 8, 7, 5, 19, 1, 14]. By equipping base stations (BSs) with an order of magnitude more antennas than conventional massive MIMO, XL-MIMO enables significantly higher spatial resolution and spectral efficiency. However, the larger array size, coupled with shrinking cell coverage in modern communication networks, increases the likelihood of user equipments operating in the near-field region. In this scenario, the conventional planar wave assumption becomes inaccurate, necessitating a shift to a more precise near-field spherical wave model. Unfortunately, the shift from planar to spherical wave propagation invalidates many fundamental techniques used in current MIMO systems, including beam training, channel estimation, and beamforming. In particular, the near-field spherical wave model poses significant challenges for beam training and channel estimation, which are critical for realizing the promised gains in spectral efficiency. Thus, investigating their impact and developing effective solutions is essential for the successful deployment of XL-MIMO.
@@
The channel information obtained from beam training is inherently limited, as only the signal strength is fed back. To fully exploit the potential of XL-MIMO and enhance multi-user communications, comprehensive channel estimation is necessary to acquire the complete channel state information (CSI). Conventional channel estimation techniques employed in massive MIMO, such as minimum mean square error (MMSE) and least squares (LS), can be directly extended to time division duplex (TDD)-based XL-MIMO systems with fully digital precoding. However, for TDD systems with hybrid precoding—commonly adopted in XL-MIMO to minimize the number of radio-frequency (RF) chains—and frequency division duplex (FDD) systems, significant challenges arise when applying existing compressive sensing (CS)-based algorithms developed for conventional massive MIMO. These methods typically exploit channel sparsity [^1] To avoid potential ambiguity, it should be emphasized that the term channel sparsity in the context of near-field XL-MIMO refers to sparsity in the parameter domain rather than structural sparsity of the channel matrix itself. in the angular domain to reduce pilot overhead, an assumption that no longer holds under the near-field spherical wave propagation model. Hence, novel channel estimation algorithms that jointly leverage sparsity in both the angular and distance domains are essential for effective CSI acquisition in near-field XL-MIMO systems.
@@
While several recent surveys have explored near-field MIMO systems [15, 8, 7, 5, 19, 1], our work is distinguished by its focused scope and systematic depth. Unlike [15, 8, 7], which provide broad overviews of near-field MIMO across multiple domains, this article concentrates specifically on the core challenges of beam training and channel estimation—two essential pillars for unlocking the potential of XL-MIMO in near-field environments. Compared to the more concise tutorial-style magazine papers [5, 19, 1], our survey offers a more structured and comprehensive taxonomy of state-of-the-art techniques tailored to these two problems. In particular, [5] and [19] focus mainly on channel estimation for localization applications, while [1] emphasizes the role of the effective beam-focusing Fraunhofer distance (EBFD) in beam training. By contrast, our work not only synthesizes recent advances in both beam training and channel estimation but also highlights open research directions, offering a holistic and in-depth resource for advancing near-field XL-MIMO research.
This article is organized as follows: Section II introduces the fundamentals of XL-MIMO. Sections III and IV present the key techniques for beam training and channel estimation in XL-MIMO systems, respectively. Section V highlights potential challenges and open directions, while Section VI draws the conclusion.
Figure 1: Spherical wave illustration of near-field channel.
Figure 2: Beam energy-spread effect (received beam pattern under far-field beamformers for different user angles and ranges under 256 antennas at 30 GHz).
@@ II Fundamentals of XL-MIMO Systems
XL-MIMO has garnered significant attention as a key enabler for 6G and beyond networks. By deploying an order of magnitude more antennas than conventional massive MIMO at the BSs, XL-MIMO not only enhances multi-user connectivity but also substantially improves spectral and energy efficiency. Notably, studies have demonstrated that XL-MIMO can achieve a considerable capacity gain over existing MIMO systems. However, the dramatic increase in array size leads to an extended Rayleigh/Fraunhofer distance, a fundamental criterion for delineating the near-field and far-field regions. Concurrently, the continuous reduction in cell size in modern communication networks exacerbates this effect, making it increasingly unlikely for user equipment to remain within the conventional far-field region. In the near-field regime, the conventional planar wave assumption becomes invalid, necessitating the adoption of a more accurate spherical wave-based channel model (shown in Fig. 1 to capture the spatial characteristics of signal propagation.
@@
In a spherical wave-based channel model, the phase of the array response vector is no longer a linear function of the array element indices, as observed in the conventional planar wave assumption. Instead, it is intricately coupled with the element-wise propagation distance in a nonlinear manner. This nonlinear phase-distance relationship significantly alters the spatial characteristics of wavefront propagation in XL-MIMO systems. As illustrated in Fig. 2, this phenomenon gives rise to the so-called beam energy-spread effect when applying conventional DFT-based beamforming [2]. Specifically, in the far-field regime, DFT beams are designed under the assumption of a linear phase progression across array elements, ensuring coherent beamforming gain at the intended direction. However, the phase evolution is nonlinear in the near-field region such that the direct application of DFT beams leads to energy dispersion over a wider angular span, reducing beamforming gain and impairing both beam training and channel estimation accuracy. This degradation poses a fundamental challenge in XL-MIMO system design.
Figure 3: Finite depth of beamfocusing (512 antennas at 100 GHz).
@@
On the other hand, the inherent coupling of phase and distance in the array response vector enables finite-depth beam focusing, a unique property of near-field spherical wavefronts, as shown in Fig. 3. Unlike far-field beams, which form a plane wavefront extending infinitely, near-field beams can concentrate energy at a specific spatial location, not only in angular dimensions but also in range. This capability allows BSs to distinguish multiple users that share the same angular direction but are located at different distances—a functionality absent in conventional MIMO systems, where angular separation is the sole basis for spatial multiplexing. The ability to focus beams at precise locations holds great potential for improving user separation, enhancing spatial multiplexing gains, and enabling new applications such as location-aware communications and holographic radio.
@@ III-A Motivation
Figure 4: Polar-domain-based beam training [2].
Although the DFT-based codebook is highly effective for beam training in far-field systems, it is not directly applicable to near-field scenarios. When a DFT-based codebook is used in the near-field, it results in beam energy dispersion, as illustrated in Fig. 2, which complicates the identification of the optimal beam angle. Additionally, applying such codewords for beamforming introduces significant power leakage to users with similar angular positions. Furthermore, the DFT-based codebook does not leverage the finite-depth beam focusing capability of the near-field, rendering it incapable of distinguishing users with identical angular coordinates but different radial distances. Therefore, for near-field XL-MIMO systems, novel beam training techniques and codebooks that jointly consider both angular and range dimensions must be developed.
Figure 5: Effective Rayleigh distance [1, 3].
@@ III-B1 Polar-domain-based beam training
Polar-domain-based beam training has emerged as a promising approach for near-field XL-MIMO systems [2]. As illustrated in Fig. 4, this method partitions the entire near-field region into discrete grids, each characterized by distinct angular and range coordinates. The codebook is then designed such that each codeword generates a beam that concentrates the majority of its energy within a single grid, effectively incorporating both angular and range information into the beam training process. Moreover, to minimize power leakage into adjacent beams, uniform sampling should be applied to the angular dimension, while non-uniform sampling is required for the range dimension to better capture the characteristics of near-field propagation.
@@
Figure 6: Beam training [16]: a) far-field for angle estimation.
Recently, several approaches have been proposed to reduce the overhead of on-grid polar-domain-based beam training [16, 24, 3]. As illustrated in Figs 6 and 7, hierarchical codebooks have been introduced to mitigate the need for a two-dimensional exhaustive search. These codebooks first perform coarse angular alignment and then refine both angle and range estimation. For instance, the authors in [16] propose a two-stage training protocol that separately handles angle and distance estimation in a uniform planar array-based system. In the first stage, a far-field codebook is used to obtain an approximate user angle. To address beam energy spread and reduce the search space in the angular domain, a wide codeword is constructed by summing multiple phase-shifted narrow far-field codewords. Once the optimal wide codeword is identified, an exhaustive search is conducted over the corresponding narrow codewords and range samples to jointly determine the user’s precise angle and range. This method achieves substantial reduction in training overhead compared to conventional beam sweeping. Additionally, a sparse linear array configuration can be employed to simultaneously generate multiple beams by leveraging near-field grating lobes. In this approach, only a subset of antennas is activated, forming a sparse array that produces a primary main lobe alongside several grating lobes [24]. As depicted in Figs 8 and 9, the corresponding multi-beam training procedure consists of two stages. First, multi-beam sweeping is performed over time to identify the best candidate beam. Next, single-beam sweeping is conducted using the entire antenna array to resolve ambiguities and pinpoint the user’s precise location. Another effective strategy to reduce training overhead is to redefine the near-field region [3]. The study in [1, 3] argues that the classical Rayleigh distance overestimates the near-field range when assessing channel capacity. Instead, the effective Rayleigh distance, which accounts for beamforming gain loss when using far-field beamforming vectors, provides a more accurate characterization of the near-field region. Under a 3 dB beamforming gain loss criterion, the effective Rayleigh distance is approximately ten times shorter than the classical definition, as shown in Fig. 5. By adopting this refined metric, the authors in [1, 3] achieve a $50\%$ reduction in codebook size compared to [2].
Figure 7: Beam training [16]: b) joint angle and distance estimation.
@@ III-B2 Conventional DFT-based beam training
Due to the beam energy spread, conventional DFT-based codebooks cannot be directly applied for angle estimation in the near-field. However, recent studies have demonstrated that the received beam pattern from a conventional DFT-based codebook can still be leveraged for user angle estimation [22, 23, 17]. In [22], a two-phase near-field beam training method was proposed to sequentially estimate the user’s angle and range, thereby mitigating the need for a computationally expensive two-dimensional exhaustive search. Specifically, the conventional DFT-based codebook is initially employed for beam sweeping. As illustrated in Fig. 2, the true user angle is approximately centered within the angular support—defined as the region where the normalized beam gain exceeds a given threshold. Once the user angle is determined, the user range is estimated via a one-dimensional exhaustive search utilizing the polar-domain codebook. Notably, the beam sweeping in [22] employs the entire antenna array, which can lead to excessive training overhead, especially in XL-MIMO systems where the number of antennas is substantial.
@@
It is important to note that both [22] and [23] utilize the polar-domain codebook for range estimation. As an on-grid beam training method, it only optimally aligns with users whose range coordinates coincide with predefined sampling points. However, since actual user positions are continuous and randomly distributed, any misalignment between the user’s true position and the nearest sampling grid results in beam training loss, thereby degrading overall system performance. A straightforward approach to improve beam training accuracy is to increase the density of range sampling points. However, this comes at the cost of increased training overhead. To address this limitation, an off-grid beam training method was proposed in [17], which eliminates the need for a polar-domain codebook by relying solely on a conventional DFT-based codebook for joint angle and range estimation. In this method, the user angle is first identified using the DFT-based codebook, as in [22]. The user range is then inferred based on the observation that the angular support width at a given user angle decreases as the user range increases, as depicted in Fig. 2 (blue and red solid curves). Since the exact functional relationship between angular support width and user range is difficult to derive analytically, a surrogate function is employed to estimate the user range in a semi-closed form.
Figure 8: Beam training [24]: a) multi-beam training.
@@ IV-A Motivation
Figure 9: Beam training [24]: b) remove angle ambiguity.
Figure 10: Systems with hybrid-field channels.
@@ V-A Real Measurement Data-Based Performance Evaluation
In most existing studies, the performance of beam training and channel estimation is evaluated using artificially generated channels based on the near-field channel model. However, in practical scenarios, real-world channel characteristics may deviate from those predicted by the assumed near-field model due to factors such as hardware impairments, environmental variations, and multipath effects. Consequently, the effectiveness of the proposed beam training and channel estimation techniques under real-world conditions remains uncertain. To bridge this gap, it is crucial to assess their performance using real-world measured channel data, ensuring their robustness and practicality in actual deployment scenarios [12].
@@ V-C Beam Training and Channel Estimation for FR3 Band
Frequency Range 3 (FR3), spanning approximately 7–24 GHz, is anticipated to be one of the most critical new spectrum bands for 6G. Its propagation characteristics differ significantly from those observed in both mmWave)/THz bands and sub-6 GHz frequencies. However, research on beam training and channel estimation for the FR3 band remains in its early stages, requiring substantial investigation. A key challenge in this context is determining whether FR3 channels will primarily exhibit near-field characteristics or if they will operate in a hybrid regime, incorporating both near-field and far-field components. Additionally, given the broad frequency range of FR3, its lower and upper portions may exhibit distinct propagation behaviors, further complicating channel modeling and system design. Therefore, a deeper understanding of FR3 channel spatial properties and their impact on beam training and channel estimation is essential for optimizing future 6G communication systems [18].
V-D Machine Learning and AI-Assisted Beam Training and Channel Estimation
Another promising research direction is the integration of machine learning and artificial intelligence (AI) techniques into beam training and channel estimation for XL-MIMO systems. Traditional approaches rely heavily on analytical channel models and handcrafted algorithms, which may not fully capture the complex propagation characteristics in near-field environments. Data-driven methods can learn channel representations and beamforming strategies directly from observations, enabling more efficient beam prediction, channel parameter estimation, and codebook design. By leveraging additional side information such as user location, sensing data, or historical channel measurements, learning-based approaches may further reduce training overhead and improve estimation accuracy. However, important challenges remain, including dataset requirements, generalization across environments, and the design of lightweight models suitable for practical deployment.
@@ VI Conclusion
XL-MIMO is poised to play a pivotal role in next-generation wireless communication systems, offering unprecedented improvements in spectral efficiency and spatial resolution. However, the transition from conventional massive MIMO to XL-MIMO introduces new challenges, particularly in beam training and channel estimation, due to the shift from planar to spherical wave propagation. In this article, we provided a structured and in-depth review of the state-of-the-art beam training and channel estimation methods, categorizing existing approaches and analyzing their effectiveness in near-field propagation environments. Looking ahead, future research efforts should focus on developing adaptive and scalable solutions, potentially leveraging sensing-aided techniques to enhance system performance. Moreover, experimental validation using real-world channel measurements will be essential to ensure practical applicability. In particular, integrating machine learning techniques with model-based signal processing may provide powerful tools for improving beam training efficiency and channel estimation accuracy in near-field XL-MIMO systems.